# DGAE: Diffusion-Guided Autoencoder for Efficient Latent Representation Learning

### v1

@@ Abstract.
Autoencoders empower state-of-the-art image and video generative models by compressing pixels into a latent space through visual tokenization. Although recent advances have alleviated the performance degradation of autoencoders under high compression ratios, addressing the training instability caused by GAN remains an open challenge. While improving spatial compression, we also aim to minimize the latent space dimensionality, enabling more efficient and compact representations. To tackle these challenges, we focus on improving the decoder’s expressiveness. Concretely, we propose DGAE, which employs a diffusion model to guide the decoder in recovering informative signals that are not fully decoded from the latent representation. With this design, DGAE effectively mitigates the performance degradation under high spatial compression rates. At the same time, DGAE achieves state-of-the-art performance with a 2×2$\times$2 × smaller latent space. When integrated with Diffusion Models, DGAE demonstrates competitive performance on image generation for ImageNet-1K and shows that this compact latent representation facilitates faster convergence of the diffusion model.
@@ 1. Introduction
The trade-off between compression and reconstruction fidelity remains a non-trivial problem to resolve. DCAE (Chen et al., 2024a) addresses this challenge by incorporating residual connections during both downsampling and upsampling stages, thereby achieving higher spatial compression rates without degrading reconstruction quality. In this work, we investigate a complementary direction that focuses on the training objectives. As shown in fig. 1(a), we observe that the commonly used GAN loss (Larsen et al., 2016) not only enhances the perceptual quality of reconstructed images, but that increasing the discriminator capacity further mitigates the degradation typically caused by aggressive compression. We hypothesize that a stronger discriminator provides richer learning signals, thus enhancing the expressiveness of the decoder.
@@ 2.1. VAEs
where $q_{\phi}(z|x)$ is the variational posterior approximated by the encoder, and $p_{\theta}(x|z)$ is the generative likelihood modeled by the decoder. The first term, also referred to as the reconstruction term, depends on the assumed form of $p_{\theta}(x|z)$—for example, under the common assumption that $p_{\theta}(x|z)$ is a Gaussian distribution with fixed unit variance, this term becomes equivalent to a mean squared error ($\ell_{2}$) loss. Consequently, the overall VAE training objective comprises the reconstruction loss $\mathcal{L}_{\text{REC}}(X,\hat{X})$ and the KL divergence loss $\mathcal{L}_{\text{KL}}$.
@@ 4. Experiments
DGAE-B 157M 128 512 122M 4.91 24.60 0.74
DGAE-M 310M 192 768 276M 4.45 25.32 0.76
DGAE-L 525M 256 1024 491M 4.40 25.56 0.77
@@ 5. Related Work
Diffusion models. Diffusion models  (Song et al., 2021; Vahdat et al., 2021; Ho et al., 2020; Nichol and Dhariwal, 2021; Karras et al., 2022; Peebles and Xie, 2023; Ma et al., 2024a; Kingma et al., 2021) have supplanted traditional generative models, including GANs  (Goodfellow et al., 2014; Karras et al., 2021; He et al., 2019, 2021) and VAEs  (Kingma and Welling, 2014; Dai and Wipf, 2019; Higgins et al., 2017; Rezende and Viola, 2018), emerging as the predominant framework in the realm of visual generation. Due to direct optimization in the pixel space and the use of multi-timestep training and inference, diffusion models were initially applied only to the synthesis of low-resolution visual content (Vahdat et al., 2021; Ho et al., 2020; Nichol and Dhariwal, 2021; Karras et al., 2022). To scale them up to high-resolution image generation, subsequent works either adopt super-resolution techniques to increase the generated images to higher resolutions  (Ramesh et al., 2022; Saharia et al., 2022; Ho et al., 2022b) or perform optimization in the latent space instead of the pixel space (Rombach et al., 2022). In parallel, another line of research focuses on accelerating the sampling process through methods such as knowledge distillation  (Salimans and Ho, 2022; Sauer et al., 2023; Song et al., 2023) and noise scheduling  (Kingma et al., 2021; Nichol and Dhariwal, 2021; Kong and Ping, 2021). By adopting these strategies, diffusion models have achieved remarkable results in visual generation (Ho et al., 2022a; Yang et al., 2024; Podell et al., 2024; Blattmann et al., 2023a; Polyak et al., 2024; NVIDIA et al., 2025; Saharia et al., 2022; Dong et al., 2023; Ma et al., 2025; Lab and etc., 2024; Zheng et al., 2024; Chen et al., 2024c; Kahatapitiya et al., 2024; Chen et al., 2024e, d, b; Xiao et al., 2024; Xie et al., 2024a; Peng et al., 2025b; Ma et al., 2024b; Chen et al., 2024f; Bar-Tal et al., 2024; Bao et al., 2024; Qin et al., 2024; Xing et al., 2024; Zhang et al., 2023; Xu et al., 2024; Liu et al., 2025; Zhou et al., 2025; Xie et al., 2024b; Tan et al., 2024; Peng et al., 2025a).
Visual Autoencoders. Due to the success of LDM’s SD-VAE, substantial efforts have been devoted to developing better autoencoders. To enable a more efficient denoising process, follow-up works have focused on improving reconstruction accuracy under high spatial compression (Chen et al., 2024a; Esser et al., 2024; Dai et al., 2023; Tian et al., 2024a; HaCohen et al., 2024).
Another popular trend is employ wavelet transforms to enhance high-frequency details  (Lab and etc., 2024; NVIDIA et al., 2025; Yu et al., 2025).In addition to the continuous AEs explored in this work, multiple discrete AEs (Van Den Oord et al., 2017; Razavi et al., 2019; Wang et al., 2024; NVIDIA et al., 2025; Tang et al., 2024; Yu et al., 2024; Li et al., 2024; Zhao et al., 2024b; Tian et al., 2024b) are proposed to aid autoregressive tasks (Sun et al., 2024; Han et al., 2024; Esser et al., 2021).

### v2

@@ Abstract.
Autoencoders empower state-of-the-art image and video generative models by compressing pixels into a latent space through visual tokenization. Although recent advances have alleviated the performance degradation of autoencoders under high compression ratios, addressing the training instability caused by GAN remains an open challenge. While improving spatial compression, we also aim to minimize the latent space dimensionality, enabling more efficient and compact representations. To tackle these challenges, we focus on improving the decoder’s expressiveness. Concretely, we propose DGAE, which employs a diffusion model to guide the decoder in recovering informative signals that are not fully decoded from the latent representation. With this design, DGAE effectively mitigates the performance degradation under high spatial compression rates. At the same time, DGAE achieves state-of-the-art performance with a $2\times$ smaller latent space. When integrated with Diffusion Models, DGAE demonstrates competitive performance on image generation for ImageNet-1K and shows that this compact latent representation facilitates faster convergence of the diffusion model.
@@ 1. Introduction
The trade-off between compression and reconstruction fidelity remains a non-trivial problem to resolve. DCAE (Chen et al., 2024f) addresses this challenge by incorporating residual connections during both downsampling and upsampling stages, thereby achieving higher spatial compression rates without degrading reconstruction quality. In this work, we investigate a complementary direction that focuses on the training objectives. As shown in fig. 1(a), we observe that the commonly used GAN loss (Larsen et al., 2016) not only enhances the perceptual quality of reconstructed images, but that increasing the discriminator capacity further mitigates the degradation typically caused by aggressive compression. We hypothesize that a stronger discriminator provides richer learning signals, thus enhancing the expressiveness of the decoder.
@@ 2.1. VAEs
[…]
@@ 4. Experiments
DGAE-B 157M 128 512 122M 4.23 25.73 0.76
DGAE-M 310M 192 768 276M 3.76 26.10 0.76
DGAE-L 525M 256 1024 491M 3.58 26.13 0.77
@@ 5. Related Work
Diffusion models. Diffusion models  (Song et al., 2021; Vahdat et al., 2021; Ho et al., 2020; Nichol and Dhariwal, 2021; Karras et al., 2022; Peebles and Xie, 2023; Ma et al., 2024a; Kingma et al., 2021) have supplanted traditional generative models, including GANs  (Goodfellow et al., 2014; Karras et al., 2021; He et al., 2019, 2021) and VAEs  (Kingma and Welling, 2014; Dai and Wipf, 2019; Higgins et al., 2017; Rezende and Viola, 2018), emerging as the predominant framework in the realm of visual generation. Due to direct optimization in the pixel space and the use of multi-timestep training and inference, diffusion models were initially applied only to the synthesis of low-resolution visual content (Vahdat et al., 2021; Ho et al., 2020; Nichol and Dhariwal, 2021; Karras et al., 2022). To scale them up to high-resolution image generation, subsequent works either adopt super-resolution techniques to increase the generated images to higher resolutions  (Ramesh et al., 2022; Saharia et al., 2022; Ho et al., 2022b) or perform optimization in the latent space instead of the pixel space (Rombach et al., 2022). In parallel, another line of research focuses on accelerating the sampling process through methods such as knowledge distillation  (Salimans and Ho, 2022; Sauer et al., 2023; Song et al., 2023) and noise scheduling  (Kingma et al., 2021; Nichol and Dhariwal, 2021; Kong and Ping, 2021). By adopting these strategies, diffusion models have achieved remarkable results in visual generation (Ho et al., 2022a; Yang et al., 2024; Podell et al., 2024; Blattmann et al., 2023a; Polyak et al., 2024; NVIDIA et al., 2025; Saharia et al., 2022; Dong et al., 2023; Ma et al., 2025; Lab and etc., 2024; Zheng et al., 2024; Chen et al., 2024a; Kahatapitiya et al., 2024; Chen et al., 2024e, d, c; Xiao et al., 2024; Xie et al., 2024a; Peng et al., 2025a; Ma et al., 2024b; Chen et al., 2024b; Bar-Tal et al., 2024; Bao et al., 2024; Qin et al., 2024; Xing et al., 2024; Zhang et al., 2023; Xu et al., 2024; Liu et al., 2025; Zhou et al., 2025; Xie et al., 2024b; Tan et al., 2024; Peng et al., 2025b).
Visual Autoencoders. Due to the success of LDM’s SD-VAE, substantial efforts have been devoted to developing better autoencoders. To enable a more efficient denoising process, follow-up works have focused on improving reconstruction accuracy under high spatial compression (Chen et al., 2024f; Esser et al., 2024; Dai et al., 2023; Tian et al., 2024b; HaCohen et al., 2024).
Another popular trend is employ wavelet transforms to enhance high-frequency details  (Lab and etc., 2024; NVIDIA et al., 2025; Yu et al., 2025).In addition to the continuous AEs explored in this work, multiple discrete AEs (Van Den Oord et al., 2017; Razavi et al., 2019; Wang et al., 2024; NVIDIA et al., 2025; Tang et al., 2024; Yu et al., 2024; Li et al., 2024; Zhao et al., 2024b; Tian et al., 2024a) are proposed to aid autoregressive tasks (Sun et al., 2024; Han et al., 2024; Esser et al., 2021).